OpenStudy (anonymous):

Find the inverse of: h(x) = 5(x-9)^(3)

jimthompson5910 (jim_thompson5910):

\[\Large h(x) = 5(x-9)^{3}\] \[\Large y = 5(x-9)^{3}\] \[\Large x = 5(y-9)^{3}\] \[\Large \frac{x}{5} = (y-9)^{3}\] \[\Large \sqrt[3]{\frac{x}{5}} = y-9\] \[\Large \sqrt[3]{\frac{x}{5}}+9 = y\] \[\Large y=\sqrt[3]{\frac{x}{5}}+9\] So the inverse function is \[\Large h^{-1}(x)=\sqrt[3]{\frac{x}{5}}+9\]